A financial analyst has been following Biostar Inc., a new high-growth company. She
estimates that the current risk-free rate is 6.25 percent, the market risk premium is 5
percent, and that Biostar Inc beta is 1.75. The current earnings per share (EPS0) is $2.50.
The company has a 40 percent payout ratio. The analyst estimates that the company's
dividend will grow at a rate of 25 percent this year, 20 percent next year, and 15 percent
the following year. After three years the dividend is expected to grow at a constant rate of
7 percent a year. The company is expected to maintain its current payout ratio. The analyst
believes that the stock is fairly priced. What is the current price of the stock?
1
Expert's answer
2016-02-20T00:00:58-0500
Use the SML equation to solve for rs
rs = 0.0625 + (0.05)*(1.75) = 0.15 = 15%
Calculate dividend per share, D0:
(EPS0)*(Payout ratio) = D0
($2.50)*(0.4) = $1.00
Calculate the dividend and price stream (once the stock becomes a constant growth stock):
D0 = $1.00; D1 = $1.00*1.25 = $1.25; D2 = $1.25*1.20 = $1.50; D3 = $1.50*1.15 = $1.725; D4 = $1.725*1.07 = $1.8458.
Use the cash flow register to calculate PV, and then solve for NPV = $18.53.
So, the answer is $18.53.